OpenStudy (anonymous):

Find the extrema of 2sinx + cos2x within the interval [0, 2pi]. Please show detailed steps. Thank you : )

OpenStudy (anonymous):

First you need to compute the derivative.

OpenStudy (anonymous):

as Thomas9 said, we find first f'(x). We know that we'll have maximum and minimum points where f'(x)=0 so we'll have extremas at points where f'(x)=0. f(x)=2sinx+cos2x f'(x)=2cosx-2sin2x then we set f'(x)=0 and find the values of x for which f'(x)=0 f'(x)=0 2cosx-2sin2x=0 (2cosx-2sin2x)/2=0/2 cosx-sin2x=0 recall from trigonometric double angle identities that sin2x=2sinxcosx so we'll have cosx-(2sinxcosx)=0 cosx-2sinxcosx=0 cosx(1-2sinx)=0 cosx=0 1-2sinx=0 x=cos^(-1)0 -2sinx=-1 x=90 degrees sinx=-1/-2 x=pi/2 sinx=1/2 x=sin^(-1)1/2 x=30 degrees x=pi/6 so the points where the function will have extremas are the points where x=pi/2 and x=pi/6

OpenStudy (anonymous):

The part I'm confused about is how you got from cosx - 2sinxcosx = 0 to cosx(1-2sinx)

OpenStudy (anonymous):

So it's cos(x)-2sin(x)cos(x)=0 Now you can just factor: cos(x)(1-2sin(x))=0 Just like: x-2xy=x(1-2y)

OpenStudy (anonymous):

@johnwessolom88 actually its an algebraic property so if you have ab-ac=a(b-c) here, both terms have cosx so we can gladly take it out to have cosx-2sincosx=cosx(1-2sinx)

OpenStudy (anonymous):

is just reversing distribution property

OpenStudy (anonymous):

just like that

OpenStudy (anonymous):

***Facepalm*** duh Sorry